# A Lightweight Framework for GPU-Accelerated Robot Learning github.com/mujocolab/mjlab

### v1

@@ 2  Design Philosophy and Scope
mjlab provides depth and raycast sensors for geometric perception. High-fidelity RGB rendering is out of scope. This does not preclude vision-based policies: a common approach is to train privileged policies using full state, then distill into vision-based controllers using external rendering.
@@ 3  Architecture
Entity 1(G1 robot)MJCF + configsEntity 2(Cube)MJCFEntity 3(Terrain)PythonMjSpecMjModelMuJoCo Warp(GPU, NN worlds)ManagerBasedRlEnvObsActionRewardTermEventCmdCurricRSL-RLTrained Policy
@@
MuJoCo Warp. MuJoCo Warp Google DeepMind and NVIDIA (2025) (docs) is a GPU-accelerated backend for MuJoCo built on NVIDIA Warp (Macklin, 2022). It preserves MuJoCo’s (Todorov et al., 2012) familiar MjModel/MjData paradigm: MjModel holds the static kinematic and dynamic description of the scene, while MjData carries the time-varying simulation state. The key addition is a leading world dimension: a single MjData object holds the state of $N$ independent simulation instances in parallel, enabling thousands of environments to be stepped in parallel. Model parameters are shared across all worlds by default, and individual fields can be expanded to vary per-world when domain randomization requires it. mjlab further captures the simulation step as a CUDA graph: the kernel execution sequence is recorded once and replayed on subsequent calls, eliminating CPU-side dispatch overhead.
@@
Manager-based paradigm. On top of the simulation layer, mjlab adopts the manager-based environment design introduced by Isaac Lab. Users define their environment by composing small, self-contained terms—reward functions, observation computations, domain randomization events—and register them with the appropriate manager. Each manager handles the lifecycle of its terms: calling them at the right point in the simulation loop, aggregating their outputs, and exposing diagnostics. Terms can be plain functions for stateless computations or classes that inherit from ManagerTermBase when they need to cache expensive setup (e.g., resolving regex patterns to joint indices at initialization) or maintain state across steps with a per-episode reset() hook. This abstraction introduces overhead compared to writing monolithic step functions, but benefits in modularity, testability, and rapid iteration outweigh the cost—a lesson reinforced by our experience with the more minimal MuJoCo Playground API, where the absence of such structure made environments harder to maintain and extend.
@@ 4  Components
Sensors. Sensors in mjlab follow a layered design. Users can wrap sensors already defined in their MJCF files and read them through MuJoCo’s native sensordata buffer, or define them entirely in Python configuration. On top of this native layer, mjlab provides custom sensor types that go beyond what MuJoCo offers: a ray-cast sensor for terrain height scanning, a contact sensor that tracks collision forces between body pairs with air-time and force-history bookkeeping, and an experimental tiled-rendering camera. All sensors implement a common interface with per-step caching and optional debug visualization.
Actuators. Actuators follow the same layered philosophy. Users can wrap actuators defined in MJCF—preserving the exact behavior specified in the model file—or create MuJoCo-native actuator types (motor, position, velocity, muscle) from configuration. Beyond these native options, mjlab provides custom implementations that compute torques on the GPU outside of MuJoCo: an ideal PD controller, a DC motor model with velocity-dependent torque saturation, and a learned MLP actuator for capturing hardware-specific dynamics from data. All three tiers coexist within the same entity. Actuation delay—common in real robots—is modeled by a wrapper class that buffers control signals and replays them with a latency quantized to the physics timestep.
Terrain. The terrain module generates ground geometry for locomotion training. Users can spawn a flat ground plane or compose a grid of sub-terrain patches, each drawn from a library of terrain types. Two families are provided: primitive terrains built from MuJoCo box geoms (flat surfaces, pyramid stairs, random-height grids) and heightfield terrains for smoother, continuous profiles (sloped pyramids, uniform noise, sinusoidal waves). Every terrain type accepts a difficulty parameter that scales its features—step height, slope angle, noise amplitude—from trivial to challenging. The grid generator supports two modes: random, where each patch is sampled independently, and curriculum, where rows encode increasing difficulty so that the curriculum manager can advance robots to harder terrain as performance improves. Users can subclass the base terrain configuration to define custom terrain types (Figure 3).
@@ 5  Manager-Based API
Action manager. The action manager receives the policy’s output tensor, splits it according to the registered action terms, and routes each slice to the appropriate entity’s actuators. It supports action clipping and tracks the two most recent actions, which observation and reward terms can reference (e.g., to compute an action rate penalty).
@@
Event manager. Events are hooks that execute at specific points in the environment lifecycle: at startup, on reset, or at fixed intervals during training. The most common use case is domain randomization. When an event modifies a field of the MuJoCo model (e.g., friction coefficients or link masses), mjlab expands that field from a shared value to a per-world array and rebuilds the CUDA capture graph to reflect the new memory layout. This happens transparently to the user, who simply specifies properties and ranges to randomize over.

### v2

@@ 2  Design Philosophy and Scope
mjlab provides ray-cast, depth, and RGB camera sensors for geometric and visual perception. Photorealistic rendering is out of scope. This does not preclude vision-based policies: a common approach is to train privileged policies using full state, then distill into vision-based controllers using external rendering.
@@ 3  Architecture
Entity 1(G1 robot)MJCF + configsEntity 2(Cube)MJCFEntity 3(Terrain)PythonMjSpecMjModelMuJoCo Warp(GPU, NN worlds)ManagerBasedRlEnvObsActionRewardTermEventCmdCurricMetricsRSL-RLTrained Policy
@@
MuJoCo Warp. MuJoCo Warp (Google DeepMind and NVIDIA, 2025) (docs) is a GPU-accelerated backend for MuJoCo built on NVIDIA Warp (Macklin, 2022). It preserves MuJoCo’s (Todorov et al., 2012) familiar MjModel/MjData paradigm: MjModel holds the static kinematic and dynamic description of the scene, while MjData carries the time-varying simulation state. The key addition is a leading world dimension: a single MjData object holds the state of $N$ independent simulation instances in parallel, enabling thousands of environments to be stepped in parallel. Model parameters are shared across all worlds by default, and individual fields can be expanded to vary per-world when domain randomization requires it. mjlab further captures the simulation step as a CUDA graph: the kernel execution sequence is recorded once and replayed on subsequent calls, eliminating CPU-side dispatch overhead.
@@
Manager-based paradigm. On top of the simulation layer, mjlab adopts the manager-based environment design introduced by Isaac Lab. Users define their environment by composing small, self-contained terms—reward functions, observation computations, domain randomization events—and register them with the appropriate manager. Each manager handles the lifecycle of its terms: calling them at the right point in the simulation loop, aggregating their outputs, and exposing diagnostics. Terms can be plain functions for stateless computations or classes that inherit from ManagerTermBase when they need to cache expensive setup (e.g., resolving regex patterns to joint indices at initialization) or maintain state across steps with a per-episode reset() hook. This abstraction introduces overhead compared to writing monolithic step functions, but benefits in modularity, testability, and rapid iteration outweigh the cost, a lesson reinforced by our experience with the more minimal MuJoCo Playground API, where the absence of such structure made environments harder to maintain and extend.
@@ 4  Components
Sensors. Sensors in mjlab follow a layered design. Users can wrap sensors already defined in their MJCF files and read them through MuJoCo’s native sensordata buffer, or define them entirely in Python configuration. On top of this native layer, mjlab provides custom sensor types that go beyond what MuJoCo offers: a BVH-accelerated ray-cast sensor that supports grid and pinhole ray patterns and returns per-ray distances, hit positions, and surface normals; a contact sensor that tracks collision forces between body pairs with air-time and force-history bookkeeping; and a GPU-accelerated camera sensor that renders RGB and depth across all worlds in parallel and supports both perspective and orthographic projection. All sensors implement a common interface with per-step caching and optional debug visualization.
Actuators. Actuators follow the same layered philosophy. Users can wrap actuators defined in MJCF, preserving the exact behavior specified in the model file, or create MuJoCo-native actuator types (motor, position, velocity, muscle) from configuration. Beyond these native options, mjlab provides custom implementations that compute torques on the GPU outside of MuJoCo: an ideal PD controller, a DC motor model with velocity-dependent torque saturation, and a learned MLP actuator for capturing hardware-specific dynamics from data. All three tiers coexist within the same entity. Actuation delay, common in real robots, is modeled by a wrapper class that buffers control signals and replays them with a latency quantized to the physics timestep.
Terrain. The terrain module generates ground geometry for locomotion training. Users can spawn a flat ground plane or compose a grid of sub-terrain patches, each drawn from a library of terrain types. Two families are provided: primitive terrains built from MuJoCo box geoms (flat surfaces, stairs, stepping stones, narrow beams, and random-height grids) and heightfield terrains for smoother, continuous profiles (sloped pyramids, Perlin noise, and sinusoidal waves). Every terrain type accepts a difficulty parameter that scales its features—step height, slope angle, noise amplitude—from trivial to challenging. The grid generator supports two modes: random, where each patch is sampled independently, and curriculum, where rows encode increasing difficulty so that the curriculum manager can advance robots to harder terrain as performance improves. Users can subclass the base terrain configuration to define custom terrain types (Figure 3).
@@ 5  Manager-Based API
Action manager. The action manager receives the policy’s output tensor, splits it according to the registered action terms, and routes each slice to the appropriate entity’s actuators. It supports action clipping and tracks the two most recent actions, which observation and reward terms can reference (e.g., to compute an action rate penalty). Shipped action terms include joint-space position and velocity targets as well as a differential inverse kinematics term that converts task-space pose commands into joint displacements via damped least-squares, with optional null-space posture regularization and soft joint-limit avoidance.
@@
Event manager. Events are hooks that execute at specific points in the environment lifecycle: at startup, on reset, or at fixed intervals during training. The most common use case is domain randomization. mjlab provides a domain randomization module covering body, joint, geometry, camera, light, material, and tendon parameters, each with configurable operations (absolute, additive, multiplicative) and sampling distributions (uniform, log-uniform, Gaussian). When an event modifies a field of the MuJoCo model, mjlab expands that field from a shared value to a per-world array and rebuilds the CUDA capture graph to reflect the new memory layout. For inertial parameters, the module implements physically consistent randomization based on the pseudo-inertia parameterization of Rucker and Wensing (2022), guaranteeing that randomized mass, center of mass, and rotational inertia remain fully physically consistent for any perturbation magnitude. The breadth of randomizable parameters, from link geometry and inertial properties to joint dynamics and actuator gains, combined with physical consistency, makes the module applicable beyond sim-to-real transfer to problems such as system identification and hardware co-design.
@@
Metrics manager. The metrics manager computes and logs custom per-episode metrics that are not tied to any specific manager, such as task success rates or tracking errors, providing a unified interface for monitoring training progress.